# TAPTRv3: Spatial and Temporal Context Foster Robust Tracking of Any Point in Long Video

### v1

@@ 4.2 Implementation Detail
Following previous methods [26, 25], we use ResNet-50 as the backbone to extract image features of every frame, 2 encoder layers with deformable attention [55] to further enhance the image features. Benefiting from our improvement, only 4 decoder layers are required to achieve optimal performance. For the supervision of location and visibility prediction, we utilize the L1 loss and binary cross-entropy loss respectively as in previous works. We use the AdamW [28] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and set the weight decay to $1\times 10^{-4}$. We train TAPTRv3 on a cluster of 8 NVIDIA A100 GPUs with a batch size of 8 in total. To make the training process more stable, we accumulate gradients 4 times to approximate a batch size of 32. After the training of TAPTRv3, we freeze it and add the additional global matching for the second stage of training. Since there are only a few parameters to be trained in this stage, it only requires about 5,300 iterations to converge.
For efficiency, in ablation studies in Sec. 4.4, we have a few modifications in our experimental settings. We reduce the number of encoder and decoder layers to 1 and 3 respectively, resize the resolution of the input video to $384\times 384$, and also reduce the number of tracking points on each video to 200. The ablation studies are conducted on 4 GeForce RTX3090 GPUs for about 28,000 iterations with randomly sampled half-size training and evaluation sets.
@@ 4.3 Comparison with the State of the Arts
We evaluate TAPTRv3 on the TAP-Vid [6] and RoboTAP [43] benchmarks and compare TAPTRv3 with previous methods to demonstrate the superiority of our approach. As shown in Table 1, TAPTRv3 achieves state-of-the-art on most metrics on these four datasets. In particular, on the Kinetics, RGB-Stacking, and RoboTAP subsets, which are composed of long videos, TAPTRv3 shows a large improvement (9.3 AJ on average) compared to TAPTRv2 [25] and surpasses the previous SoTA method significantly by about 2.4 AJ on average. On the DAVIS dataset with shorter video, although it is slightly behind TAPTRv2 on AJ, there is still an improvement in $<\delta_{avg}^{x}$, indicating better localization capability.
Although Cotracker3 [19] and BootsTAPIR [8] achieve remarkable performance, they both introduced extra internal real-world data for training. Specifically, CoTracker3 re-renders the Kubric training set with a length of 64 frames, which narrows the gap in video length between training and evaluation. Then an additional 15K real-world videos are introduced for fine-tuning. BootsTAPIR trains their model on the original Kubric training set but introduces an extra 15M real-world video clips, which is approximately 1,360 times more than the synthetic data (11K) we use for training. The results show that despite these methods utilizing much more additional data for training, TAPTRv3 still achieves competitive performance.
Note that, for fair comparison, we do not utilize global matching to help reestablish tracking here. When the global matching is also considered, the results are further improved. For more detail please refer to Sec. 4.4.
Figure 5: Visualization of ground truth, TAPTRv3, and TAPTRv2. In the early stages of tracking, both TAPTRv2 and TAPTRv3 can produce accurate tracking results. As the camera shifts and the tracking points disappear from the video for an extended period (over 100 frames), TAPTRv2 completely loses the tracking, whereas TAPTRv3 maintains stable tracking. Best view in electronic version.
@@ 4.4 Ablation Studies and Analysis
We start our ablation from TAPTRv2 and conduct ablation studies on every key component in TAPTRv3 to validate their effectiveness. We further conduct some more detailed ablations to investigate the best implementation choice. To focus on the ability of TAPTRv3 in handling long videos, we conduct ablations on Kinetics.
@@
Visibility-aware Long-Temporal Attention. We first replace the RNN-like long-temporal modeling with the long-temporal attention. As shown in Table 4, the comparison between Row 2 and Row 1 shows that the replacement provides a significant improvement (2.9 AJ), indicating the superiority of the attention mechanism over RNN in handling varying data length, which aligns with the findings in modern LLMs. Furthermore, the comparison between Row 3 and Row 2 shows that enabling the long-temporal attention to utilize visibility prediction to reduce noises caused by occlusion will further improve the performance by 1.2 AJ. The significant improvements brought by the VLTA module indicate the effectiveness of incorporating richer temporal information.
@@
Discarding Sliding Window. With VLTA, the model captures temporal information from all previous frames, making it redundant to recompute temporal attention within a small window. Therefore, we eliminate the sliding window by reducing the window size from 8 to 1. In this case, the original temporal attention in TAPTRv2 will degenerate to an MLP with residual. We retain this module in the following ablations for fair comparison. As shown in Table 4, the comparison between Row 4 and Row 3 shows a 1.0 AJ improvement. We attribute this to better position initialization. More specifically, initializing the position of the target tracking point in the current frame with the estimation from a more nearby frame simplifies the estimation process, as the distance that the model needs to refine is significantly reduced.
@@
Context-aware Cross-Attention. As shown in Table 4, the comparison between Row 6 and Row 5 shows that the introduction of CCA further brings a significant improvement of 1.0 AJ. The improvement verifies the effectiveness of CCA in improving the robustness of the spatial feature querying.
@@
Ignore the Position Supervision of Invisible Points. It is worth noting that, predicting the position of the target tracking point when it is occluded is an ill-posed problem. Forcing the model to localize the occluded point can destabilize the learning process and may lead the model to learn a bias toward a fixed motion pattern. As shown in Rows 5 and 4 of Table 4, simply ignoring the supervision of invisible points’ location predictions results in an improvement of 0.8 AJ.
@@
Patch-level Similarity Calculation in CCA. As shown in Table 4, we conduct comparative experiments on different methods to obtain patch-level similarity. The results show that the “Element-wise”, which only considers the similarities between two points that are located at the same position in the two patches, is less effective than the “Every two point” strategy that is adopted in our current CCA. We believe this may be because the “Every two point” strategy has an advantage in handling more complex spatial variations, such as rotation, and is more tolerant of the sampling point’s location. For more details, please refer to our supplementary.
@@
Spatial Context Updating. As shown in Tabel 4, neither utilizing our VLTA nor an MLP to update the query’s context features yields better results. This indicates that maintaining the target tracking points’ spatial context throughout the tracking process not only reduces computation costs but also helps spatial feature querying. For more details about this ablation please refer to supplementary.
Ablation on Tracking Reestablish with Global Matching As discussed in Sec. 3.4, TAPTRv3 triggers global matching only when a scene cut is detected. This is because we find that, as shown in Table 5, solely relying on global matching to provide the initial location of the point queries in every frame obtains inferior performance. To quantitatively demonstrate the tracking improvements brought by global matching, we conducted comparative experiments on the Kinetics dataset and its subset, which is composed of videos with scene cuts, accounting for about 27% of the full set. As shown in Table 6, Although the gain from global matching is not substantial on the full set, it brings an average improvement of 0.7 AJ on videos with scene cuts, which is significant. This demonstrates the effectiveness of the global matching module and the mechanism we use to trigger it.
Input Positions of Decoder AJ $<\delta^{x}_{avg}$ OA
Global Matching Calculation 51.0 63.3 84.1
Previous Frame’s Prediction 52.9 65.8 87.2
Table 5: Ablation on decoder input positions.
Input Positions of Decoder Dataset AJ $<\delta^{x}_{avg}$ OA
Previous Frame’s Prediction Kin. 54.5 67.5 88.2
Global Matching if S.C. Kin. 54.6 67.8 88.4
Previous Frame’s Prediction Kin. w/ S.C. 54.4 66.9 86.6
Global Matching if S.C. Kin. w/ S.C. 55.1 68.1 87.4
Table 6: Ablation on global matching trigger. Whether to employ the global matching positions when detecting scene cuts. “Kin.” is short for TAP-Vid-Kinetics, and “S.C.” indicates scene cuts.
@@ 5 Visualization
We visualize the tracking results of TAPTRv2 and TAPTRv3 in a long video (about 250 frames) in Fig. 5. Note that, for a fair comparison, we do not add global matching. As the camera shifts away from the player and pans around for a long period (approximately 100 frames), TAPTRv2 loses track completely. Whereas, TAPTRv3 maintains stable tracking even after the target has disappeared for a long duration without the help of global matching, demonstrating the robustness of TAPTRv3 in tracking any point in long videos. For more visualizations, please refer to our supplementary.
@@ 6 Conclusion
In this paper, we have presented TAPTRv3, a strong method for the TAP task. TAPTRv3 improves TAPTRv2 primarily by developing the Context-aware Cross-Attention (CCA) and Visibility-aware Long-Temporal Attention (VLTA) to address the shortage of feature querying. CCA utilizes the spatial context to help obtain more robust attention weights in cross-attention, leading to a better perception of 2D image space. VLTA replace the RNN-like long-temporal modeling with an attention mechanism to enable the perception of longer temporal context while avoiding the feature drifting issue. VLTA further utilizes the visibilities to improve the quality of long-temporal attention, leading to a better feature querying ability along the temporal dimension. Additionally, TAPTRv3 further improves its performance in long videos by introducing the global matching module with a carefully designed trigger strategy. With the help of these novel designs, TAPTRv3 surpasses TAPTRv2 by a large margin and obtains state-of-the-art performance on multiple challenging datasets. Even when compared with the methods trained on extra internal data, TAPTRv3 remains competitive.
Supplementary Material
Appendix A Appendix
In this appendix, we provide more ablation studies on the design of our key components and some hyper-parameters (See Sec. A.1). We then provide more details about the implementation and experimental settings (see Sec.A.2). Finally, we present some more visualization results to show the superiority of TAPTRv3 and the effectiveness of our designs (See Sec. A.3 and the supplementary videos).
@@ A.1 More Ablation Studies
[…]
Number of Context Features $N^{2}$ AJ $<\delta^{x}_{avg}$ OA
$N^{2}=1$ 51.9 64.8 86.9
$N^{2}=9$ 52.9 65.8 87.2
$N^{2}=25$ 52.1 65.3 87.4
@@
Memory Size of VLTA. TAPTRv3 eliminates the sliding window [26, 25] and introduces the Visibility-aware Long-Temporal Attention (VLTA) module to extend the temporal attention to an arbitrary length, while considering the visibility. Excluding the use of visibility, the recent SAM2 [56]777[56]Nikhila Ravi, Valentin Gabeur, Yuan-Ting Hu, Ronghang Hu, Chaitanya Ryali, Tengyu Ma, Haitham Khedr, Roman R¨adle, Chloe Rolland, Laura Gustafson, et al. SAM 2: Segment Any-thing in Images and Videos. arXiv preprint arXiv:2408.00714, 2024. 12. adopts a similar temporal modeling approach, where SAM2 maintains a memory to record features from past frames in a FIFO manner and limits the memory size to 8 to focus on recent frames. To demonstrate the advantage of extending temporal attention to arbitrary lengths and the generalization ability of our VLTA to temporal lengths, we conduct the ablation studies on the memory size. As illustrated in Table 8, the significant positive correlation between memory size and model performance indicates that, although the perception range of our VLTA is limited to 1 to 23 frames during training  [^8] Because our training data are videos with fixed lengths of 24 frames, our VLTA is still able to generalize beyond 24 frames. This generalization ability enables us to expand the range of temporal perception, allowing the collection of long-term temporal information from all past frames to help improve the robustness of long-term point tracking.
Memory Size of VLTA AJ $<\delta^{x}_{avg}$ OA
12 48.0 65.7 83.6
24 51.4 66.7 86.5
48 53.7 67.3 87.9
All Past 54.5 67.5 88.2
Table 8: Ablation on memory size of VLTA.
@@
The Method of Obtaining Similarity Map in Global Matching. In Sec. 3.4, we propose to conduct the global matching to reestablish point tracking when a scene cut is encountered. Instead of simply relying on a point-level feature to construct a similarity map in global matching as in previous works [5, 7], we propose incorporating spatial context features, similar to the CCA module, to enhance accuracy. (For more implementation details, please refer to Sec. A.2). To investigate the effectiveness of this approach, we conduct an ablation study on it. As shown in Table 9, the incorporation of spatial context features brings a relatively small performance improvement. However, considering that we only trigger global matching when a scene cut is detected and its computational cost is negligible, we adopt the use of spatial context features in TAPTRv3.
Feature of Tracking Point AJ $<\delta^{x}_{avg}$ OA
Point-level Feature 55.07 67.96 87.16
Spatial Context Features 55.10 68.06 87.37
Table 9: Ablation on calculation of similarity map
@@
To avoid misunderstanding, we emphasize that our main contribution in global matching lies in the novel triggering mechanism, rather than the global matching itself.
A.2 More Implementation Details

### v2

@@ 4.2 Implementation Detail
Unlike previous works (Li et al., 2024c; b), we use Resnet-18 instead of Resnet-50 as the backbone for higher efficiency. In the Transformer, we employ two encoder layers with deformable attention (Zhu et al., 2021) to further enhance the image features. While benefiting from our improvement, only 4 decoder layers are required to achieve optimal performance. For the supervision of location and visibility prediction, we utilize the L1 loss and binary cross-entropy loss, respectively, as in previous works. We use the AdamW (Loshchilov, 2017) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and set the weight decay to $1\times 10^{-4}$. We train TAPTRv3 on a cluster of 8 NVIDIA A100 GPUs for about 33,000 iterations with a batch size of 8 in total. To make the training process more stable, we accumulate gradients 4 times to approximate a batch size of 32. After the training of TAPTRv3, we freeze it and add the additional global matching for the second stage of training. Since there are only a few parameters to be trained in this stage, it only requires about 5,300 iterations to converge.
For efficiency, in ablation studies in Sec. 4.4, we have a few modifications in our experimental settings. We reduce the number of encoder and decoder layers to 1 and 3, respectively, resize the resolution of the input video to $384\times 384$, and also reduce the number of tracking points on each video to 200. The ablation studies are conducted on 4 GeForce RTX3090 GPUs for about 33,000 iterations with randomly sampled half-size training and evaluation sets.
@@ 4.3 Comparison with the State of the Arts
We evaluate TAPTRv3 on the Kinetics, RGB-Stacking, and RoboTAP datasets, which have relatively long videos, and compare it with previous methods to demonstrate its superiority on long videos. As shown in Table 1, TAPTRv3 achieves state-of-the-art on most metrics on these three datasets. With our insights and careful designs, TAPTRv3 shows a significant improvement (9.2 AJ on average) compared to TAPTRv2 (Li et al., 2024b) even with a more lightweight backbone and fewer decoder layers. Meanwhile, compared with the previous state-of-the-art Track-On (Aydemir et al., 2025) that uses DINOv2 (Oquab et al., 2023) as the backbone, we achieve an average improvement of 1.0 AJ.
Although Cotracker3 (Karaev et al., 2024a) and BootsTAPIR (Doersch et al., 2024) achieve remarkable performance, they both introduced extra internal real-world data for training. Specifically, CoTracker3 re-renders the Kubric training set with a length of 64 frames, which narrows the gap in video length between training and evaluation. Then, an additional 15K real-world videos are introduced for fine-tuning. BootsTAPIR trains its model on the original Kubric training set but introduces an extra 15M real-world video clips, which is approximately 1,360 times more than the synthetic data (11K) we use for training. The results show that despite these methods utilizing much more additional data for training, TAPTRv3 still achieves competitive performance. Note that, for fair comparison, we do not utilize auto-triggered global matching to help reestablish tracking here. When the global matching is also enabled, the results are further improved, as shown in Table 6 of Sec. 4.4.
Table 2: Ablations on key component of TAPTRv3. “LTA” refers to Long-Temporal Attention, “Vis-Aware” is short for visibility-aware, “Re. Win.” indicates the removal of the sliding window, and “Sup. Vis.” represents only supervising the visible points’ positions during training.
Row LTA Vis-Aware Re. Win. Sup. Vis. CCA AJ
1 44.5
2 47.8
3 48.8
4 49.5
5 51.1
6 52.9
Table 3: Ablation on patch-level similarity calculation.
Patch-level Similarity AJ $\delta^{x}_{avg}$ OA
Element-wise 51.3 64.8 87.4
Every two point 52.9 65.9 87.8
Table 4: Ablation of methods for context features updating.
Update Methods AJ $\delta^{x}_{avg}$ OA
VLTA 51.2 64.4 86.1
MLP 51.7 65.4 87.7
No Updates 52.9 65.9 87.8
@@ 4.4 Ablation Studies and Analysis
We start our ablation from TAPTRv2 and conduct ablation studies on every key component in TAPTRv3 to validate their effectiveness. We further conduct some more detailed ablations to investigate the best implementation choice. To focus on the ability of TAPTRv3 in handling long videos, we conduct ablations on TAP-Vid-Kinetics.
@@
Visibility-aware Long-Temporal Attention. We first replace the RNN-like long-temporal modeling with the long-temporal attention. As shown in Table 4, the comparison between Row 2 and Row 1 shows that the replacement provides a large improvement (3.3 AJ), indicating the superiority of the attention mechanism over RNN in handling varying length, which aligns with the findings in modern LLMs. Furthermore, the comparison between Row 3 and Row 2 shows that enabling the long-temporal attention to utilize visibility prediction to reduce noises caused by occlusion will further improve the performance by 1.0 AJ. The significant improvements brought by the VLTA validate the effectiveness of incorporating richer temporal information.
@@
Removal of the Sliding Window. With VLTA, the model captures temporal information from previous frames, making it redundant to recompute temporal attention within a small window. Therefore, we eliminate the sliding window by reducing the window size from 8 to 1. In this case, the original temporal attention in TAPTRv2 will degenerate to an MLP with residual connections. We retain this module in the following ablations for fair comparison. As shown in Table 4, the comparison between Row 4 and Row 3 shows a 0.7 AJ improvement. We attribute this to better position initialization. More specifically, initializing the position of the target tracking point in the current frame with the estimation from a nearby frame simplifies the estimation process. In addition, this modification also enables TAPTRv3 to process input videos in a streaming manner.
@@
Context-aware Cross-Attention. As shown in Table 4, the comparison between Row 6 and Row 5 shows that the introduction of CCA further brings a significant improvement of 1.8 AJ, verifying the effectiveness of CCA in improving the robustness of the spatial feature querying.
@@
Only Supervise the Position of Visible Points. It is worth noting that predicting the position of the target tracking point when it is occluded is an ill-posed problem. Forcing the model to localize the occluded point can destabilize the learning process and may lead the model to learn a bias toward a fixed motion pattern. As shown in Rows 5 and 4 of Table 4, simply ignoring the supervision of invisible points’ location predictions results in an improvement of 1.6 AJ.
@@
Patch-level Similarity Calculation in CCA. As shown in Table 4, we conduct comparative experiments on different methods to obtain patch-level similarity. The results show that the “Element-wise”, which only considers the similarities between two points that are located at the same position in the two patches, is less effective than the “Every two point” strategy that is adopted in our current CCA. This is because the “Every two point” strategy has an advantage in handling more complex spatial variations, such as rotation, and is more tolerant of the sampling point’s location. For more details, please refer to Sec. A.4 in our appendices.
@@
Spatial Context Updating. As shown in Table 4, neither utilizing our VLTA nor an MLP to update the query’s context features yields better results. This indicates that maintaining the target tracking points’ spatial context throughout the tracking process not only reduces computation costs but also helps spatial feature querying. For more details, please refer to Sec. A.4 in our technical appendix.
Ablation on Auto-Triggered Global Matching As discussed in Sec. 3.4, TAPTRv3 triggers global matching only when a scene cut is detected. This design is based on the empirical finding that naively applying global matching at every frame results in inferior performance, as shown in Table 6. Therefore, TAPTRv3 defaults to using the previous frame’s prediction for initialization. It activates global matching only when a scene cut is detected, which serves to re-establish tracking and prevent failure. This automatic trigger mechanism boosts the performance of our best model by 0.2 AJ, as shown in Table 6. We further construct a subset of Kinetics by selecting all videos with scene cuts. On this subset, this module brings an improvement of 0.5 AJ in the same experiment setting, confirming its effectiveness. More details can be found in Sec. A.1 of the appendix.
Table 5: Ablation on the input positions of decoder.
Input Positions of Decoder AJ $\delta^{x}_{avg}$ OA
Global Matching Calculation 51.1 64.0 86.5
Previous Frame’s Prediction 52.9 65.9 87.8
Table 6: Ablation on global matching. Whether to use auto-triggered global matching.
Input Positions of Decoder Dataset AJ $\delta^{x}_{avg}$ OA
Previous Frame’s Prediction Kin. 54.9 67.5 88.2
Global Matching if S.C. Kin. 55.1 67.7 88.4
@@ 5 Visualization
We visualize a qualitative comparison between TAPTRv2 and TAPTRv3 on a challenging video sequence of approximately 250 frames in Fig. 3. The camera pans left until the person and goal are out of view, remains so for about 100 frames, and then pans back. During this long-term absence, TAPTRv2 completely loses the initial target points. In contrast, TAPTRv3 maintains stable tracking throughout the sequence. This result highlights the effectiveness of our proposed components in mitigating feature drifting and achieving more accurate localization. Notably, for fair comparison, the global matching mechanism in TAPTRv3 was disabled.
Figure 3: Visual comparison between TAPTRv3 and TAPTRv2.
@@ 6 Conclusion
In this paper, we have presented TAPTRv3, a strong method for the TAP task. TAPTRv3 improves TAPTRv2 primarily by developing the Context-aware Cross-Attention (CCA) and Visibility-aware Long-Temporal Attention (VLTA) to address the shortage of feature querying. CCA improves key-aware deformable attention by leveraging spatial context, which helps point-level tasks obtain more robust and accurate attention weights for updating both features and positions. VLTA replaces the RNN-like long-temporal modeling with an attention mechanism, enabling the perception of longer temporal context and mitigating the issue of feature drifting. VLTA further utilizes the visibilities to improve the quality of long-temporal attention, leading to a better feature querying ability along the temporal dimension. Additionally, TAPTRv3 further improves its performance in long videos by introducing the auto-triggered global matching mechanism. With the help of our insights and these novel designs, TAPTRv3 surpasses TAPTRv2 by a large margin and obtains state-of-the-art performance on multiple challenging datasets. Even when compared with the methods trained on extra internal data, TAPTRv3 remains competitive.
Appendix A Technical Appendices and Supplementary Material
In this appendix, we provide more ablation studies on the design of our key components and some hyper-parameters (see Sec. A.1). Beyond demonstrating the superior performance of our model, we also analyze its efficiency, including inference speed and GPU memory usage (see Sec. A.2). In addition to showcasing our model’s superior performance on long videos, we also provide a comparison with other methods on the DAVIS dataset in TAP-Vid (Doersch et al., 2022) benchmark, which consists of relatively shorter videos (see Sec.A.3). Furthermore, we provide more details on the implementation and experimental settings (see Sec.A.4). In addition, we also discuss the limitations of our model to provide a comprehensive understanding (see Sec.A.5). Finally, we present some more visualization results to show the superiority of TAPTRv3 and the effectiveness of our designs (see Sec. A.6 and the supplementary videos).
@@ A.1 More Ablation Studies
Number of Context Features. As discussed in Sec. 3.1 and Sec. 3.3, we sample $N^{2}$ features in the form of a grid with 1-pixel grid interval when preparing the point query’s context features $\mathbf{C}$ and each sampling point’s context features $\mathbf{K}_{t}^{m}$ in cross-attention. At different scales of the feature map, the grid interval is the same. This allows us to fuse information from receptive fields of different sizes. As $N$ increases, it effectively means that the patch size becomes larger, and the context features tend to represent more global information. To find the optimal value of $N$, we conduct experiments with N=1N=1444When $N=1$, the CCA will actually degenerate to the vanilla key-aware deformable attention, we can also find this performance in Row 5 of Table 4., $N=3$, and $N=5$ using the same settings as Sec. 4.4. The results are shown in Table 7, where $N=3$ obtains the best performance. Compared with sampling 25 context features ($N=5$), sampling 9 context features ($N=3$) around the point not only requires fewer computing resources but also yields better performance, with an AJ improvement of 0.7. These results indicate that for the task of point tracking, context features are important, but they should not be so excessive or represent too large areas of the image as this could prevent accurate description of the tracked points.
@@
Memory Size of VLTA. TAPTRv3 eliminates the sliding window (Li et al., 2024c; b) and introduces the Visibility-aware Long-Temporal Attention (VLTA) module to extend the temporal attention to an arbitrary length, while considering the visibility. Excluding the use of visibility, the recent SAM2  [^5] Nikhila Ravi, Valentin Gabeur, Yuan-Ting Hu, Ronghang Hu, Chaitanya Ryali, Tengyu Ma, Haitham Khedr, Roman R¨adle, Chloe Rolland, Laura Gustafson, et al. SAM 2: Segment Any-thing in Images and Videos. arXiv preprint arXiv:2408.00714, 2024. 12. adopts a similar temporal modeling approach, where SAM2 maintains a memory to record features from past frames in a FIFO manner and limits the memory size to 8 to focus on recent frames. To demonstrate the advantage of extending temporal attention to arbitrary lengths and the generalization ability of our VLTA to temporal lengths, we conduct ablation studies on the memory size. Specifically, we use the trained TAPTRv3 model from Sec 4.3 and perform evaluations on the Kinetics dataset with different memory sizes. As illustrated in Table 8, the significant positive correlation between memory size and model performance indicates that, although the perception range of our VLTA is limited to 1 to 23 frames during training  [^6] Because our training data are videos with fixed lengths of 24 frames, our VLTA is still able to generalize beyond 24 frames. This generalization ability enables us to expand the range of temporal perception, allowing the collection of long-term temporal information from all past frames to help improve the robustness of long-term point tracking.
Table 8: Ablation on the memory size of VLTA (Kinetics).
Memory Size of VLTA AJ $\delta^{x}_{avg}$ OA
12 51.9 65.2 86.5
24 53.1 66.3 87.3
48 54.5 67.2 88.0
All Past 54.9 67.5 88.2
However, unconstrained temporal memory size will lead to CUDA out-of-memory problems in practical applications. We find that when processing videos with up to 3,500 frames, the GPU memory requirement reaches 24GB because of the large amount of cached temporal memory. To balance performance and efficiency, and to enable the model to handle videos of arbitrary length, we need to limit the size of the temporal memory and manage it with FIFO mechanism. To this end, we further evaluated RoboTAP (with videos up to 1,300 frames) with more diverse memory sizes. As shown in Fig. 4, the performance converges when the memory size reaches 512. With the help of temporal memory management, TAPTRv3 is capable of handling downstream applications, which usually require the model to process online streaming videos.
Figure 4: Ablation on the memory size of VLTA (RoboTAP).
@@
The Method of Obtaining Similarity Map in Global Matching. In Sec. 3.4, we propose the auto-triggered global matching to reestablish point tracking when a scene cut is encountered to prevent the loss of tracking targets. Instead of simply relying on a point-level feature to construct a similarity map in global matching as in previous works (Cho et al., 2024; Doersch et al., 2023), we propose incorporating spatial context features, similar to the CCA module, to enhance accuracy. (For more implementation details, please refer to Sec. A.4). To investigate the effectiveness of this approach, we conduct an ablation study on it. As shown in Table 9, the incorporation of spatial context features brings a relatively small performance improvement. However, considering that we only trigger global matching when a scene cut is detected and its computational cost is negligible, we adopt the use of spatial context features in TAPTRv3.
Table 9: Ablation on calculation of similarity map.
Feature of Tracking Point AJ $\delta^{x}_{avg}$ OA
Point-level Feature 55.75 67.88 87.57
Spatial Context Features 55.79 67.94 87.57
@@
To avoid misunderstanding, we emphasize that our main contribution in global matching lies in the novel auto-trigger mechanism, rather than the global matching itself.
More Ablation on Auto-Triggered Global Matching. We evaluate the proposed auto-triggered global matching mechanism on the full Kinetics dataset. As shown in Table 6, while the method yields performance gains, the improvement is relatively modest. This is largely because only a subset of videos in Kinetics contain scene cuts, and thus actually activate the mechanism, whereas evaluation metrics are averaged over the entire dataset. To better isolate the effectiveness of our approach, we construct a scene-cut subset comprising all videos in Kinetics that contain at least one scene cut, accounting for approximately 27% of the 1,144 total videos. We conduct the same ablation study on this subset. As shown in Table 10, triggering global matching upon detecting scene cuts leads to a significant improvement (0.5 AJ), demonstrating the effectiveness of our method in relevant scenarios.
Table 10: Ablation on auto-triggered global matching. Whether to trigger the global matching when encountering scene cuts on the Kinetics scene-cut subset. “Kin.” is short for TAP-Vid-Kinetics, and “S.C.” indicates scene cuts
Input Positions of Decoder Dataset AJ $\delta^{x}_{avg}$ OA
Previous Frame’s Prediction Kin. w/ S.C. 55.3 67.0 86.9
Global Matching if S.C. Kin. w/ S.C. 55.8 67.9 87.6
A.2 Efficiency Analysis
Inference Speed. Notably, while TAPTRv3 introduces two additional modules, CCA and VLTA, which increase computational overhead compared to TAPTRv2, the inference speed of TAPTRv3 is even faster than that of TAPTRv2. We benchmark the runtime performance on the TAP-Vid-DAVIS dataset using a single GeForce RTX3090 GPU, measuring the total number of video frames and the model’s inference time to calculate the average FPS. We also include CoTracker (Karaev et al., 2024b), which is a highly representative and influential method, as a baseline. For all experiments, we set the input video resolution to $384\times 512$, with an average of 22 tracked points per video. Each experiment was repeated five times to report a stable average. For a fair comparison, we set the window size to 8 and the window stride to 4 for all models. As shown in Table 11, under the same conditions, TAPTRv3 achieves 15 FPS higher than TAPTRv2.
Table 11: Average FPS of TAPTRv2 and TAPTRv3.
Method Average FPS Method Average FPS Method Average FPS
Cotracker 26.4 TAPTRv2 41.9 TAPTRv3 57.2
We attribute the faster speed to the following reasons. (1) The CCA module introduces only 9 more sparse sample points, with the additional computational overhead negligible compared to the overall network. (2) While we introduce VLTA, we also remove the Temporal Attention module and the complex Window Post-processing in TAPTRv2. (3) TAPTRv3 uses a more lightweight backbone (Resnet-18 vs. Resnet-50) and fewer decoder layers (4 layers vs. 5 layers), making the model more efficient. Additionally, it utilizes lower-resolution training data ($384\times 512$ vs. $512\times 512$), resulting in lower-resolution inputs during inference as well.
GPU Memory Overhead. When the memory size in VLTA is set to 512 (as discussed in Sec A.1), our model is capable of processing videos of arbitrary length in a streaming manner. When tracking 100 points simultaneously in a streaming video, the GPU memory usage is less than 2GB, making its deployment cost-effective.
Table 12: Comparison of TAPTRv3 with different backbone and training resolution.
Kinetics RGB-Stacking RoboTAP DAVIS
Method AJ $\delta^{x}_{avg}$ OA AJ $\delta^{x}_{avg}$ OA AJ $\delta^{x}_{avg}$ OA AJ $\delta^{x}_{avg}$ OA
TAPTRv3 (Resnet-50, $512\times 512$) 54.5 67.5 88.2 73.0 86.2 90.0 64.6 77.2 90.1 63.2 76.7 91.0
TAPTRv3 (Resnet-18, $384\times 512$) 54.9 67.5 88.2 72.3 84.1 90.8 64.5 77.3 89.7 63.2 76.4 90.6
Table 13: Comparison of TAPTRv3 with prior methods. We use to indicate the introduction of additional training data. Specifically, CoTracker3 additionally incorporates 15K real videos. BootsTAPIR and BootsTAPNext-B train on an additional 15M real videos. Anthro-LocoTrackR leverages an extra 1.4K human motion data. TAPTRv3 obtains state-of-the-art performance on most datasets and remains competitive with methods trained on extra internal data. Training data: (Kub24), (Kub48), and (Kub64) refer to Kubric (Greff et al., 2022) with 24, 48, and 64 frames per video, respectively. (PO) PointOdyssey (Zheng et al., 2023), (FT) FlyingThings++ (Mayer et al., 2016). For fair comparison, we do not utilize auto-triggered global matching here.
Training DAVIS
Method Data AJ $\delta^{x}_{avg}$ OA
PIPs (Harley et al., 2022) FT 42.2 64.8 77.7
PIPs++ (Zheng et al., 2023) PO 69.1
TAP-Net (Doersch et al., 2022) Kub24 33.0 48.6 78.8
TAPIR (Doersch et al., 2023) Kub24 56.2 70.0 86.5
CoTracker (Karaev et al., 2024b) Kub24 61.8 76.1 88.3
TAPTR (Li et al., 2024c) Kub24 63.0 76.1 91.1
TAPTRv2 (Li et al., 2024b) Kub24 63.5 75.9 91.4
LocoTrack (Cho et al., 2024) Kub24 63.0 75.3 87.2
CoTracker3 (online) (Karaev et al., 2024a) Kub64 64.5 76.7 89.7
Track-On (online) (Aydemir et al., 2025) Kub24 65.0 78.0 90.8
BootsTAPIR (Doersch et al., 2024) Kub24+15M 61.4 73.6 88.7
CoTracker3 (online) (Karaev et al., 2024a) Kub64+15K 63.8 76.3 90.2
Anthro-LocoTrack (Kim et al., 2025) Kub24+1.4K 64.8 77.3 89.1
BootsTAPNext-B (online) (Zholus et al., 2025) Kub48+15M 65.2 78.5 91.2
TAPTRv3 (online) Kub24 63.2 76.4 90.6
A.3 More Comparisons
We compare TAPTRv3 with previous methods on the DAVIS subset of the TAP-Vid (Doersch et al., 2022) benchmark. DAVIS contains relatively short videos, with an average length of only 70 frames. As shown in Table 13, on shorter videos TAPTRv3 performs comparably to TAPTRv2. Although it is slightly behind in AJ, there is still an improvement in $\delta_{avg}^{x}$, indicating better localization capability. The result is reasonable, as our design is mainly intended to improve tracking performance on long videos, which has already been validated in Table 1 of the main paper.
A.4 More Inplementation Details
Backbone and Training Resolution. As discussed in Sec. 4.1 and Sec. 4.2, compared to TAPTR and TAPTRv2, which use Resnet-50 as the backbone to extract image features, TAPTRv3 adopts the more lightweight Resnet-18. Additionally, TAPTRv3 is trained on videos with a resolution of $384\times 512$, which is lower than the $512\times 512$ resolution used by TAPTR and TAPTRv2. This is primarily because our experiments showed that a lighter backbone and lower training resolution can achieve comparable performance, as indicated by the results in Table 12. These modifications improve the efficiency of our model and also demonstrate the superiority of the decoder design.